OpenStudy (anonymous):

What is an equation for the line with slope 2/3 and y-intercept 9?

pooja195 (pooja195):

y=mx+b m=slope b=y intercept

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

How would I answer that? Help me through it please

OpenStudy (anonymous):

Do you need answers?

pooja195 (pooja195):

Ok y=mx+b This is known as slope intercept form y intercept= b

OpenStudy (anonymous):

y=2/3x y=9x y=2/3x+9 y=9x+2/3

OpenStudy (anonymous):

Oh right

OpenStudy (anonymous):

It's hard to figure out how to do in an equation though or slope I mean

pooja195 (pooja195):

Can you plug in what i said and make th equation? y=mx+b m=slope=2/3 b=y intercept =9

OpenStudy (anonymous):

No, litteraly I don't understand what you're giving me here lol

pooja195 (pooja195):

y=mx+b <--slope intercept form Replace b with y intercept which is 9 and replace m with 2/3

OpenStudy (anonymous):

y=2/3x+9?

pooja195 (pooja195):

Well done ;)

OpenStudy (anonymous):

Is that right, If so that was easy peasy... I feel happy about my self now lol

OpenStudy (anonymous):

Thanks bro

pooja195 (pooja195):

=)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)